OpenStudy (katherinesmith):

What is the exact value of cos (pi/2) as found on the unit circle?

OpenStudy (katherinesmith):

\[\cos\frac{ \pi }{ 2 }\]

OpenStudy (anonymous):

\(cos\left(\frac{\pi}{2}\right)=0\)